OpenStudy (anonymous):

Can someone verify if this simple algebra question on my test is right or wrong?

OpenStudy (anonymous):

See attached

OpenStudy (anonymous):

I translated this as G=2W-12 and G+W=128 resulting in 128-W=2W-12

OpenStudy (amistre64):

70 is correct 128 = 2A-12 12+128 = 2A 140 = 2A 70 = A

OpenStudy (anonymous):

But it says COMBINED weight

OpenStudy (amistre64):

why they use "w" for "Adam" is beyond me....

OpenStudy (anonymous):

there formula, to me, implies 128 is just the weight of george

OpenStudy (amistre64):

Yes, combined weight is George and Adam. And you figured out that equation correctly the first time :)

OpenStudy (amistre64):

let me see it again..

OpenStudy (anonymous):

for second one...should be 40..i feel

OpenStudy (anonymous):

I get 46 2/3Lbs

OpenStudy (amistre64):

G = 2A George is twice Adams weight....

OpenStudy (amistre64):

G = 2A - 12

OpenStudy (amistre64):

G + A = 128

OpenStudy (amistre64):

2A - 12 + A - 128

OpenStudy (amistre64):

3A = 128 + 12

OpenStudy (amistre64):

gotta read thru my typos lol

OpenStudy (amistre64):

3A = 140 A = 140/3 = 46 and some odd lbs

OpenStudy (amistre64):

the test is worded wrong... human error :)

OpenStudy (anonymous):

This isn't the first time ive got mis information from this school. Thats why I come here to check. Sometimes they flutter up, sometimes its me.

OpenStudy (amistre64):

You bring it to the teachers attention and they lok it over and make the correction to your grade. Perfection is an ideal, not a reality :)

OpenStudy (anonymous):

this online program didn't really come with a teacher, just literature

OpenStudy (amistre64):

there has to be a human someplace in the setup.... :)

OpenStudy (amistre64):

a "contact" button even

OpenStudy (anonymous):

the grades aren't really important, its my understanding that is I'm almost finished with it. I think I have one more lesson/test than I'll start looking for my government job :)

OpenStudy (anonymous):

so, hey, thanks for the help

OpenStudy (anonymous):

hopefully I wont have to return for the next lesson :P

OpenStudy (amistre64):

youre welcome :) good luck!